OpenStudy (anonymous):

which of the following are equal to the expression below ? check all that apply sqrt9/sqrt81 a.3/9 b.2/9 c.3 d.1/3 ?? help

OpenStudy (experimentx):

d

OpenStudy (anonymous):

3/9 and 1/3 so A and D

OpenStudy (hoblos):

sqrt9/sqrt81 = sqrt(3^2)/sqrt(9^2) = 3/9 =1/3

Directrix (directrix):

sqrt9/sqrt81 = 3/9 = 1/3 ---> Option D